OpenStudy (anonymous):

what is the shortest distance between the line y=1/2x+3 and point (-3,-3)

OpenStudy (jdoe0001):

what do you think?

OpenStudy (jdoe0001):

 so, in short you're been really asked is "what is the equation of a line that is perpendicular to y=1/2x+3 and passes through (-3,-3)" ? so since you've done slopes, you get such equation then you'd have 2 equations, one for each line thus a system of equations recall that in a system of equations, the solution is where the graphs meet or intersect so the these two lines will intersect at a point on the 1st line, the given one then you'd have 2 points the solution to the system of equations, and (-3,3) well, to find how long that is, use the distance formula \(\bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ \square }}\quad ,&{\color{blue}{ \square }})\quad &({\color{red}{ -3}}\quad ,&{\color{blue}{ 3}}) \end{array}\qquad d = \sqrt{({\color{red}{ x_2}}-{\color{red}{ x_1}})^2 + ({\color{blue}{ y_2}}-{\color{blue}{ y_1}})^2}\)